The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught. Step 1 of 2 : Suppose a sample of 782 suspected criminals is drawn. Of these people, 273 were captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.

Step 2 of 2 :

Suppose a sample of 782782 suspected criminals is drawn. Of these people, 273273 were captured. Using the data, construct the 98%98% confidence interval for the population proportion of people who are captured after appearing on the 1010 Most Wanted list. Round your answers to three decimal places.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = р = x / n = 273 / 782 = 0.3491

Z\alpha/2 = 2.326

Margin of error = E = Z\alpha / 2 * \sqrt ((р * (1 - р )) / n)

= 1.326 * (\sqrt((0.3491 * (1 - 0.3491)) / 782)

= 0.040

A 98% confidence interval for population proportion p is ,

р - E < p < р + E

0.349 - 0.040 < p < 0.349 + 0.040

0.309 < p < 0.389

(0.309 , 0.389)